OpenStudy (anonymous):

Please Help me with this one calculate dy/dx y = 7x^2 − 9x + 19 / 2x + 4

OpenStudy (anonymous):

Is this the function? \[\large y = \frac{7x^2 − 9x + 19}{ 2x + 4}\]

OpenStudy (anonymous):

yes @CliffSedge

OpenStudy (anonymous):

Familiar with the quotient rule?

OpenStudy (anonymous):

"Low d-high minus high d-low all over low-squared"

OpenStudy (anonymous):

or f'(x)g(x)-f(x)g'(x) / 2x+4

OpenStudy (anonymous):

im just stuck on the next couple steps

OpenStudy (anonymous):

For \[\large y=\frac{f(x)}{g(x)}\] \[\large y' = \frac{g(x)f'(x)-f(x)g'(x)}{(g(x))^2}\]

OpenStudy (anonymous):

you have to square the (2x+4) on the bottom.

OpenStudy (anonymous):

g(x) = 7x^2-9x+19 right?

OpenStudy (anonymous):

I'm using g(x) for the denominator, so f(x)=7x^2-9x+19 and g(x)=2x+4

OpenStudy (anonymous):

for f'(x) did you get f'(x)=14x-9. and for g'(x)= 2

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

\[\frac{ 14x^2+56x-38 }{ (2x+4)^2 }\]

OpenStudy (anonymous):

is that the correct answer?

OpenStudy (anonymous):

Double check the constant in your numerator.

OpenStudy (anonymous):

i got \[28x^2-18x+56x-36-14x^2+18x-38\] - before solving

OpenStudy (anonymous):

\[14x^2-20x+2\]

OpenStudy (anonymous):

\(28x^2−18x+56x−36−14x^2+18x−38\) \( = 28x^2−14x^2+18x−18x+56x−36−38 \)

OpenStudy (anonymous):

Combine like terms to simplify.

OpenStudy (anonymous):

got it . 14x^2+56x-74

OpenStudy (anonymous):

Right, now also notice (if you want to simplify it), that you can factor \(\large 2^2\) out of the bottom and cancel a factor of 2 to reduce the fraction.

OpenStudy (anonymous):

oh ok. thx for the help

OpenStudy (anonymous):

My pleasure.